# A battle of designs: triangular vs. L-shaped detectors and parity violation in the gravitational-wave background

### v1

@@ VI Conclusions
Across all scenarios, ET in 2L configurations are consistently more sensitive to parity violation than triangular ET configurations, with the gap most pronounced for standalone ET. Parameter estimation studies further demonstrate that network design strongly affects the ability to constrain the GWB parameters. We find that an optimal network design choice can tighten a searched parameter space and can place polarization bounds that exclude $\Pi=0$ at 95% compared to alternative detector designs. In Appendix A, we show that parity violation detection in ET alone has been ruled out; circular polarization in a third-generation network requires CE detectors.
@@ Acknowledgements.
We thank Michael Ebersold for valuable discussions on third-generation network configurations, Alberto Mariotti and Miguel Vanvlasselaer for helpful discussions. We are also thankful to Gianmassimo Tasinato and Angelo Ricciardone for useful comments. This document has an ET document number ET-0483A-25. The work of MS is partially supported by the Science and Technology Facilities Council (STFC grant ST/X000753/1). HD is supported by the Strategic Research Program High-Energy Physics of the Research Council of the Vrije Universiteit Brussel, the iBOF “Unlocking the Dark Universe with Gravitational Wave Observations: from Quantum Optics to Quantum Gravity” of the Vlaamse Interuniversitaire Raad.
@@ Appendix A Discussion on Einstein Telescope alone
We briefly assess the detection prospects of a parity violating GWB using ET alone. For this purpose, we consider baselines that differ from those adopted in the main analysis. Specifically, we introduce the notation $\Delta^{\prime}_{10}$ and $\Delta^{\prime}_{15}$ to denote closed, fully symmetric triangular ET configurations with 10 km and 15 km arms, respectively. These correspond to the idealized case of a perfectly equilateral triangle formed by three co-located interferometers, with each vertex separated by 60°. For comparison, we also include the configurations $\Delta_{10}$ and $\Delta_{15}$ already used in the main text and defined in Ebersold et al. (2024). These differ in that the three interferometers do not form an exactly closed triangle. As a result, the $\Delta_{10}$ and $\Delta_{15}$ configurations are slightly asymmetric. This asymmetry yields enhanced sensitivity to $V$-modes, since the chiral ORF is slightly larger than that of a perfectly closed triangle.
In addition to triangular configurations, we also examine two 2L configurations, denoted by (SR)0 and (SR)45. The (SR)0 setup corresponds to the optimal 2L configuration for GWB detection where the two detectors are aligned (relative orientation of 0°). By contrast, the (SR)45 represents the most pessimistic case for GWB detection, where the detectors are rotated by 4545° with respect to one another. It is important to emphasize that (SR)0 and (SR)45 differ from (SR)1 and (SR)2 considered in the main analysis, since the latter were defined relative to the optimal configuration of a global network including two CEs, whereas here the reference is ET alone. In Table 5, the latitude and longitude of each network center, the X- and Y-arm endpoints and azimuths are given.
@@
Figure 5 shows the PI curves for both the I- and V-modes across the different ET-only configurations. For comparison, we also include a parity-violating PL GWB with spectral index $\alpha=0$, to facilitate direct comparison with the corresponding SNR prospects shown in Fig. 6. The perfectly closed and symmetric triangular configurations $\Delta^{\prime}$ exhibit noticeably lower V-mode sensitivity than the slightly asymmetric $\Delta$ setups used in the main text, illustrating that even small asymmetries in ET’s geometry can enhance parity-violation detection. Nevertheless, all triangular configurations, whether symmetric or slightly asymmetric, remain significantly less sensitive to the V-modes than the 2L designs.
@@
We have again performed SNR studies for multiple spectral indices $\alpha$. For the I-mode SNRs, the ranking from most to least sensitive is: (SR)0, (SR)45, $\Delta_{15}=\Delta_{15}^{\prime}$ and $\Delta_{10}=\Delta_{10}^{\prime}$. For the V-mode SNRs, the resulting network rankings are summarized in Tables 6, confirming that 2L configurations consistently outperform triangular ET for V-mode detection. Moreover, current bounds on the parity-violated GWB Abac et al. (2025b) have ruled out the parity-violating parameter space that is detectable by both 2L and the $\Delta$ configurations, as seen in Fig.6. Consequently, if ET were constructed without CE, none of the configurations would be able to detect circularly polarized GWBs at a confident SNR level.
$\alpha$ 3,2,1-3,-2,-1 $0,\tfrac{1}{2},1,\tfrac{2}{3},2,3$

### v2

@@ VI Conclusions
Across all scenarios, $15$ km triangular ET configurations are most sensitive to parity violation, followed by 2L configurations, while $10$ km triangular ET configurations are the least sensitive. Parameter estimation studies further demonstrate that network design strongly affects the ability to constrain the GWB parameters. We find that adopting a more sensitive network configuration can tighten the searched parameter space and can place polarization bounds that exclude $\Pi=0$ at 95% compared to less sensitive detector designs. In Appendix A, we show that parity violation detection in ET alone has been ruled out; circular polarization in a third-generation network requires CE detectors.
@@ Acknowledgements.
We are grateful to Jacopo Uggeri for extensive discussions. We also thank Michael Ebersold for valuable insights on third-generation network configurations, Alberto Mariotti and Miguel Vanvlasselaer for helpful discussions, and Gianmassimo Tasinato and Angelo Ricciardone for useful comments. This document has an ET document number ET-0483A-25. The work of MS is partially supported by the Science and Technology Facilities Council (STFC grant ST/X000753/1). HD is supported by the Strategic Research Program High-Energy Physics of the Research Council of the Vrije Universiteit Brussel, the iBOF “Unlocking the Dark Universe with Gravitational Wave Observations: from Quantum Optics to Quantum Gravity” of the Vlaamse Interuniversitaire Raad.
@@ Appendix A Discussion on Einstein Telescope alone
We briefly assess the detection prospects of a parity violating GWB using ET alone, building on previous ET-only detector design and sensitivity studies Branchesi and others (2023). For this purpose, we consider baselines that differ from those adopted in the main analysis. Specifically, we introduce the notation $\Delta^{\prime}_{10}$ and $\Delta^{\prime}_{15}$ to denote closed, fully symmetric triangular ET configurations with $10$ km and $15$ km arms, respectively. These correspond to the idealized case of a perfectly equilateral triangle formed by three co-located detectors, with each vertex separated by 60°. For comparison, we also include the configurations $\Delta_{10}$ and $\Delta_{15}$ already used in the main text and defined in Ebersold et al. (2024). These differ in that the three detectors do not form an exactly closed triangle. As a result, the $\Delta_{10}$ and $\Delta_{15}$ configurations are slightly asymmetric. This asymmetry yields enhanced sensitivity to $V$-modes, since the chiral ORF is slightly larger than that of a perfectly closed triangle.
In addition to triangular configurations, we also examine two 2L configurations, denoted by (SR)0 and (SR)45. The (SR)0 setup corresponds to the optimal 2L configuration for GWB detection where the two detectors are aligned (relative orientation of 0°). By contrast, the (SR)45 represents the most pessimistic case for GWB detection, where the detectors are rotated by 4545° with respect to one another, consistent with previous studies of 2L configurations and relative orientations Branchesi and others (2023)). It is important to emphasize that (SR)0 and (SR)45 differ from (SR)1 and (SR)2 considered in the main analysis, since the latter were defined relative to the optimal configuration of a global network including two CEs, whereas here the reference is ET alone. In Table 5, the latitude and longitude of each network center, the X- and Y-arm endpoints and azimuths are given.
@@
Fig. 5 shows the PI curves for both the $I$- and $V$-modes across the different ET-only configurations. For the $I$-mode sensitivities, the trends are consistent with previous ET-only sensitivity studies Branchesi and others (2023), while here we additionally present the corresponding $V$-mode sensitivities relevant for parity-violating GWBs. For comparison, we also include a parity-violating PL GWB with spectral index $\alpha=0$, to facilitate direct comparison with the corresponding SNR prospects shown in Fig. 6. The perfectly closed and symmetric triangular configurations $\Delta^{\prime}$ exhibit noticeably lower $V$-mode sensitivity than the slightly asymmetric $\Delta$ setups used in the main text, illustrating that even small asymmetries in ET’s geometry can enhance parity-violation detection. Nevertheless, all triangular configurations, whether symmetric or slightly asymmetric, are significantly less sensitive to the $V$-modes than the 2L designs.
@@
We have again performed SNR studies for multiple spectral indices $\alpha$. For the $I$-mode SNRs, the ranking from most to least sensitive is: (SR)0, (SR)45, $\Delta_{15}=\Delta_{15}^{\prime}$ and $\Delta_{10}=\Delta_{10}^{\prime}$. For the $V$-mode SNRs, the resulting network rankings are summarized in Tables 6, confirming that 2L configurations consistently outperform triangular ET for $V$-mode detection. We note that correlated Newtonian noise between the co-located detectors of the triangular configuration Janssens and others (2024) is not included in the present study. Accounting for this contribution would further reduce the sensitivity prospects of the triangular ET designs. Moreover, current bounds on the parity-violated GWB Abac and others (2025a) have ruled out the parity-violating parameter space that is detectable by both 2L and the $\Delta$ configurations, as seen in Fig.6. Consequently, if ET were constructed without CE, none of the configurations would be able to detect circularly polarized GWBs at a confident SNR level.
Order $\alpha$ 3,2,1-3,-2,-1 $0,\tfrac{1}{2},1,\tfrac{2}{3},2,3$